# Through the Perspective of LiDAR: A Feature-Enriched and Uncertainty-Aware Annotation Pipeline for Terrestrial Point Cloud Segmentation

### v1

@@ 4 Discussion
Across all three benchmarked datasets—Mangrove3D, ForestSemantic, and Semantic3D—feature enrichment consistently improves segmentation accuracy, but the gains are not uniform across feature types. Surface normals (N3) emerge as the most consistently valuable, delivering substantial improvements in class-level IoUs for geometrically complex structures such as tree stems, roots, and small urban objects. Basic features (I.R.Z) provide more modest but still meaningful gains, particularly for classes with strong vertical separation or height-dependent geometric–radiometric patterns. Importantly, the optimal trade-off between accuracy and computational cost is achieved with compact, nine-channel configurations (e.g., IRZ_N3_CAP), which capture nearly all available discriminative power, while avoiding the redundancy and increased training time associated with larger feature stacks. This suggests that thoughtful feature selection, rather than maximal channel inclusion, is critical for efficient TLS segmentation pipelines.
@@
Several design choices render the proposed framework both scalable and adaptable. First, all features are derived exclusively from LiDAR data, ensuring applicability to TLS scans regardless of the availability or reliability of RGB imagery. Second, feature fusion is performed at the bottleneck stage, enabling straightforward substitution of different encoder–decoder backbones (e.g., CNNs or Transformers) without altering the overall pipeline. Third, the patch size is determined adaptively, allowing the model to accommodate varying input resolutions and scene extents while maintaining computational efficiency. Fourth, the framework accommodates a flexible number of input channels under a triadic grouping scheme, thereby enabling seamless integration of diverse radiometric, geometric, or statistical descriptors without architectural modification. Finally, the pseudo-scanner center concept introduced for ForestSemantic can be extended to mobile or UAS-based LiDAR platforms that cover larger areas at lower point densities, though further optimization may be required to determine ideal center placements, spherical map resolutions, and truncation distances.
@@
Limitations, future directions, and broader implications
Persistent errors in fine-scale, self-occluding structures (e.g., mid-story stems in dense canopies) and performance drops in visually heterogeneous regions indicate that enriched geometric cues and current feature stacks remain insufficient to capture the full variability of complex natural scenes. While multi-model ensembles improve accuracy and uncertainty estimation, their added computational cost may limit scalability without substantial computational resources. Future work should investigate more robust and generic preprocessing strategies, along with lighter backbone architectures that reduce redundancy, while preserving segmentation performance. Efforts also are needed to narrow the reliability gap between uncertainty maps and actual error distributions by adopting improved uncertainty estimation methods.
@@ 5 Conclusion
While sensor fusion is becoming increasingly popular, our feature-enrichment results highlight the untapped potential of LiDAR point clouds alone. Careful exploitation of LiDAR-derived features can, in some cases, surpass early (or “raw”) LiDAR–camera fusion in the spherical projection space, revealing a distinct and informative view through the perspective of LiDAR. We therefore conclude that for downstream tasks such as semantic segmentation, rigorous feature design, and preprocessing of LiDAR data are no less important than integrating multi-source sensing.
@@ Acknowledgments
The authors are grateful to Dr. Nidhal Carla Bouaynaya and Dr. Bartosz Krawczyk for their valuable insights on uncertainty evaluation and semantic segmentation methods. We thank Mr. Brett Matzke for his assistance in setting up the computing resources. We also acknowledge RIT Research Computing https://doi.org/10.34788/0s3g-qd15 for providing access to NVIDIA A100 computing resources.
@@
Fei Zhang: Conceptualization, Methodology, Software, Validation, Formal Analysis, Investigation, Data Curation, Visualization, Writing – original draft, Project administration, Supervision. Fei Zhang and Rob Chancia: Conceptualization, Experimental design. Fei Zhang: Experiment execution. Fei Zhang and Josie Clapp: Data Annotation (Mangrove3D dataset). Rob Chancia, Richard MacKenzie, and Jan van Aardt: Data Collection (Mangrove3D dataset). Rob Chancia and Amirhossein Hassanzadeh: Validation, Writing – review & editing, Software. Jan van Aardt: Writing – review & editing, Supervision. Dimah Dera: Writing – review & editing.
@@
The Mangrove3D dataset has been deposited on Zenodo (doi:10.5281/zenodo.16933584) and is currently under embargo; it will be made publicly accessible upon publication. The accompanying preprocessing, feature engineering, training, and evaluation code (including environment specification and reproducibility instructions) is available at https://fz-rit.github.io/through-the-lidars-eye/.
@@
Appendix A Supplementary Material for Mangrove3D Dataset
@@ Appendix B Supplementary Material for Stage 1
Table 6: Preprocessing steps for the feature maps shown in Figures 21 and 23.
@@
where $\mathbf{x}_{j}\in\mathbb{R}^{3}$ denotes the coordinates of the $j$-th neighbor, $\bar{\mathbf{x}}$ is the neighborhood centroid, and $N$ is the number of points in the neighborhood.
@@
Here, curvature ($\kappa$) reflects how “linear” or “planar” a neighborhood is, approaching zero for flat surfaces and increasing when points are distributed along a sharp edge or corner. Anisotropy ($A$) quantifies the dominance of one principal axis over the others, with higher values indicating elongated, line-like structures (e.g., stems or branches). Planarity ($P$) measures the degree to which points lie within a 2D plane, with higher values indicating flat surfaces (e.g., ground or large leaves). These features provide compact measures of local surface geometry and have been widely used for point-cloud classification and segmentation harshit2022geometric , drones5040104 , ijgi10030187 .
@@
  • MNF green1988transformation : a two-step transform that first whitens data using an estimated noise covariance matrix, then applies PCA on the noise-whitened data. The resulting components are ranked by signal-to-noise ratio.
  • ICA hyvarinen2000independent : a statistical technique that seeks latent components which are mutually independent and non-Gaussian, thus capturing higher-order relationships beyond variance and correlation.
@@ Appendix C Supplementary Material for Stage 2
C.1 Annotation guideline
A concise annotation protocol is provided here [details].
C.2 Loss functions
@@
Cross-Entropy loss mao2023cross :
@@
C.3 Ensemble inference and uncertainty estimation
@@
Finally, epistemic uncertainty is estimated as the mutual information lakshminarayanan2017simple , kendall2017uncertainties :

### v3

@@ 4 Discussion
Across all three benchmarked datasets—Mangrove3D, ForestSemantic, and Semantic3D—feature enrichment consistently improves segmentation accuracy, but the gains are not uniform across feature types. Surface normals (N3) emerge as the most consistently valuable, delivering substantial improvements in class-level IoUs for geometrically complex structures such as tree stems, roots, and small urban objects. Basic features (I.R.Z) provide more modest but still meaningful gains, particularly for classes with strong vertical separation or height-dependent geometric-radiometric patterns. Importantly, the optimal trade-off between accuracy and computational cost is achieved with compact, nine-channel configurations (e.g., IRZ_N3_CAP), which capture nearly all available discriminative power, while avoiding the redundancy and increased training time associated with larger feature stacks. This suggests that thoughtful feature selection, rather than maximal channel inclusion, is critical for efficient TLS segmentation pipelines.
@@
Several design choices render the proposed framework both scalable and adaptable. First, all features are derived exclusively from LiDAR data, ensuring applicability to TLS scans regardless of the availability or reliability of RGB imagery. Second, feature fusion is performed at the bottleneck stage, enabling straightforward substitution of different encoder-decoder backbones (e.g., CNNs or Transformers) without altering the overall pipeline. Third, the patch size is determined adaptively, allowing the model to accommodate varying input resolutions and scene extents while maintaining computational efficiency. Fourth, the framework accommodates a flexible number of input channels under a triadic grouping scheme, thereby enabling seamless integration of diverse radiometric, geometric, or statistical descriptors without architectural modification. Finally, the pseudo-scanner center concept introduced for ForestSemantic can be extended to mobile or UAS-based LiDAR platforms that cover larger areas at lower point densities, though further optimization may be required to determine ideal center placements, spherical map resolutions, and truncation distances.
@@
Limitations and future directions
While the proposed multi-model ensemble improves segmentation accuracy and uncertainty estimation, its increased computational cost may limit applicability in real-time or resource-constrained settings. Future work will therefore investigate more robust and generic preprocessing strategies, as well as lighter backbone architectures or knowledge distillation approaches to reduce redundancy while preserving pseudo-label quality. Additionally, although uncertainty guidance was shown to restructure the annotation workflow and reduce manual labeling scope, this study does not report quantitative measures of annotation efficiency such as time or interaction counts; future work will incorporate systematic logging of annotation processes to enable controlled, quantitative evaluation of human annotation effort. Further efforts are also needed to narrow the gap between uncertainty estimates and actual error distributions by adopting improved uncertainty estimation methods.
@@ 5 Conclusion
While sensor fusion is becoming increasingly popular, our feature-enrichment results highlight the untapped potential of LiDAR point clouds alone. Careful exploitation of LiDAR-derived features can, in some cases, surpass early (or “raw”) LiDAR-camera fusion in the spherical projection space, revealing a distinct and informative view through the perspective of LiDAR. We therefore conclude that for downstream tasks such as semantic segmentation, rigorous feature design, and preprocessing of LiDAR data are no less important than integrating multi-source sensing.
@@ Acknowledgments
The authors are grateful to Dr. Nidhal Carla Bouaynaya and Dr. Bartosz Krawczyk for their valuable insights on uncertainty evaluation and semantic segmentation methods. We thank Mr. Brett Matzke for his assistance in setting up the computing resources. We also acknowledge RIT Research Computing of Technology [2019] for providing access to NVIDIA A100 computing resources.
@@
Fei Zhang: Conceptualization, Methodology, Software, Validation, Formal Analysis, Investigation, Data Curation, Visualization, Writing - original draft, Project administration, Supervision. Fei Zhang and Rob Chancia: Conceptualization, Experimental design. Fei Zhang: Experiment execution. Fei Zhang and Josie Clapp: Data Annotation (Mangrove3D dataset). Rob Chancia, Richard MacKenzie, and Jan van Aardt: Data Collection (Mangrove3D dataset). Rob Chancia and Amirhossein Hassanzadeh: Validation, Writing - review & editing, Software. Jan van Aardt: Writing - review & editing, Supervision. Dimah Dera: Writing - review & editing.
@@
The Mangrove3D dataset has been released on Zenodo (doi:10.5281/zenodo.16933584). The accompanying preprocessing, feature engineering, training, and evaluation code (including environment specification and reproducibility instructions) is available at https://fz-rit.github.io/through-the-lidars-eye/.
@@
Appendix A Annotation Guideline for Mangrove3D Dataset
The annotation guideline for making semantic segmentation masks of spherical maps of the Mangrove3D dataset can be accessed via this link.
@@ Appendix B Supplementary Material for Stage 1
Table 9: Preprocessing steps for the feature maps shown in Figures 22 and 24.
@@
where $\mathbf{x}_{j}\in\mathbb{R}^{3}$ denotes the coordinates of the $j$-th neighbor, $\bar{\mathbf{x}}$ is the neighborhood centroid, and $N$ is the number of points in the neighborhood. For the Mangrove3D dataset, we adopted a fixed neighborhood configuration with neighbor_radius=0.06mneighbor\_radius=0.06~m and max_neighbors=50max\_neighbors=50, selected empirically to balance robustness to point density variation and sensitivity to local geometric structure.
@@
Here, curvature ($\kappa$) reflects how “linear” or “planar” a neighborhood is, approaching zero for flat surfaces and increasing when points are distributed along a sharp edge or corner. Anisotropy ($A$) quantifies the dominance of one principal axis over the others, with higher values indicating elongated, line-like structures (e.g., stems or branches). Planarity ($P$) measures the degree to which points lie within a 2D plane, with higher values indicating flat surfaces (e.g., ground or large leaves). These features provide compact measures of local surface geometry and have been widely used for point-cloud classification and segmentation Harshit et al. [2022], Duran et al. [2021], Atik et al. [2021].
@@
  • MNF Green et al. [1988]: a two-step transform that first whitens data using an estimated noise covariance matrix, then applies PCA on the noise-whitened data. The resulting components are ranked by signal-to-noise ratio.
  • ICA Hyvärinen and Oja [2000]: a statistical technique that seeks latent components which are mutually independent and non-Gaussian, thus capturing higher-order relationships beyond variance and correlation.
@@ Appendix C Supplementary Material for Stage 2
C.1 Loss functions
@@
Cross-Entropy loss Mao et al. [2023a]:
@@
C.2 Ensemble inference and uncertainty estimation
@@
Finally, epistemic uncertainty is estimated as the mutual information Lakshminarayanan et al. [2017], Kendall and Gal [2017]: